OpenStudy (this):

A baseball is thrown up into the air. The ball's height in feet is modeled by the equation h(t) = -16t2 + 32t + 10. (h = height, t = time in seconds) Find the maximum height of the ball. a) 10 feet b) 2.3 feet c) 23 feet d) 26 feet

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

First, find the time when \(\dfrac{dh}{dt}=0\), and then find the height at that time.

OpenStudy (anonymous):

it's an upside down parabola so it would be a maximum. To find it u can simplify the equation into the turning point one eg y= a(x-b) + c and the turning point is at (b,c) or you could find the x-intercepts and plus them together and then divide by two, after u have got this x value sub in the equation to find ur y value then u will have (x,y) as ur t.p.

OpenStudy (anonymous):

simplifying the equation requires completing the square.

OpenStudy (anonymous):

@JayDs - It's much quicker with calculus.

OpenStudy (anonymous):

(as well as more generally applicable...parabolas are only special cases where you can maximize in that manner)

OpenStudy (anonymous):

oh sorry, i haven't learnt calculus yet.

OpenStudy (this):

@JayDS its fine :)

OpenStudy (this):

therefore....is it A? B? C? D???

OpenStudy (anonymous):

@this - Please give the problem an attempt. We're not just going to solve it for you.

OpenStudy (mertsj):

The maximum height of the ball is the y coordinate of the vertex. The x coordinate of the vertex is : \[x=\frac{-b}{2a}\]

OpenStudy (mertsj):

Find that, plug it into the equation in place of x and find y. That will be the maximum height of the ball.

OpenStudy (this):

i am not getting it plzz help me guys?

OpenStudy (anonymous):

@this - What is \(\dfrac{dh}{dt}\)?

OpenStudy (mertsj):

26 feet. D -b/2a=/-32/-32=1 -16(1)+32(1)+10=26 Vertex is (1,26) high point is 26